OpenStudy (karabella98):

A dock worker pushes a 50 kg crate up a 1-m-high, 3-m-long ramp. Ignoring friction, how much work did he do? A. 150 J B. 490 J C. 1,470 J D. 1,960 J

OpenStudy (anonymous):

We need to figure out gravity's force acting along the ramp. That is the force needed to push the box up the ramp. Work is then just force times distance. angle = arctan(1/3) = 18.45 degrees Fg = m*g*sin(angle) = 50 * 9.8 * sin(18.45) = 155 N So we need 155 N to counter gravity acting on the box. Over a distance of d^2 = 1^2 + 3^3 = 10 --> d = 3.16 m W = 155 * 3.16 = 490 J So the answer is (B)